OpenStudy (anonymous):

y = x2 + 3x - 10 For the above quadratic equation find: a.The y-intercept. b.The x-intercepts if possible. c.The vertex. d. The line of symmetry Please help me. by explaining it step by step.

OpenStudy (anonymous):

y intercept is the value without the variable

OpenStudy (anonymous):

This is because the y axis is at x=0

OpenStudy (anonymous):

So we can find this by plugging in zero: y = x^2 + 3x - 10 y = (0)^2 + 3(0) - 10 y = -10 This is our y intercept

OpenStudy (anonymous):

You can use the quadratic equation to find the zeros of your parabola.

OpenStudy (anonymous):

This is the quadratic equation \[\frac{ -b +/-\sqrt(b^2-4ac) }{ 2a }\] This is the form the parabola has come in :\[ax^2 + bx + c \]

OpenStudy (anonymous):

The vertex can be found by converting ax^2 + bx + c into a(x - h)^2 + k. You will need to complete the square to do this.

OpenStudy (anonymous):

Your vertex will be (h, k)

OpenStudy (anonymous):

Your line of symmetry will be k. We know it will not be h because this parabola is vertical. It would be horizontal if y was squared in our given equation.

OpenStudy (anonymous):

so the x-intercepts are? i'm confused on that @swagmaster47

OpenStudy (campbell_st):

wow... that seems complicated... here is an easy method (a) y- intercept substitute x = 0 and then solve to y... that will be the intercept (b) x - intercepts... set y = 0 then solve \[x^2 + 3x - 10 = 0\] this factors... find 2 numbers that multiply to -10 and add to 3, the larger factor is negative and the smaller is positive. this will help you to find the x- intercepts. (d) the line of symmetry is half way between the x - intercepts... so add the 2 intercepts and have them (c) substitute the line of symmetry value into the equation and the y- value you get along with the x value for line of x symmetry form the vertex hope it helps

OpenStudy (campbell_st):

oops larger factor is positive and the smaller is negative in (b)